OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±6) and asymptotes at y = ±(3/2)x.

OpenStudy (anonymous):

y^2/36 - x^2/16 = 1 y^2/4 - x^2/ 9 = 1 y^2/ 36 - x^2/4 = 1 y^2/ 16 - x^2/ 36 = 1